# Self-Interacting Dark Matter with Mass Segregation: A Unified Explanation of Dwarf Cores and Small-Scale Lenses

### v2

@@ Abstract
In two-component self-interacting dark matter (SIDM) models with inter-species interactions, mass segregation arises naturally from collisional relaxation, enhancing central densities and gravothermal evolution. We demonstrate that models with velocity-dependent interactions, both within and between species, can connect several small-scale observations while remaining consistent with cluster-scale constraints. This combination enables core formation in dwarf halos, where the presence of baryons increases the inner densities and enhances the predicted strong lensing signatures. Using cosmological and controlled simulations alongside an accurate parametric model, we show that this framework explains the structure of dark perturbers observed in strong lensing systems, and significantly increases the efficiency of small-scale lenses by a factor of $\sim 5$, consistent with the galaxy-galaxy strong lensing excess reported in clusters. Importantly, mass segregation can enhance the Einstein radii of SIDM halos relative to their CDM counterparts, overcoming a key challenge in one-component SIDM scenarios. Our results establish mass segregation in two-component SIDM as a self-consistent and testable model capable of simultaneously addressing multiple small-scale challenges in structure formation.
I Introduction
@@
In this work, we demonstrate that SIDM with mass segregation provides a natural unified framework for resolving these tensions. Mass segregation in multi-component SIDM models arise as a consequence of collisions among particles of different masses. In the two-component case, inter-species scatterings lead to energy transfer from the heavier to the lighter species, resulting in denser inner halos Yang et al. (2025a); Patil and Fischer (2025). This process is negligible in CDM due to inefficient gravitational relaxation among particles (unlike stars Zhong and Shapiro (2025)), but becomes significant in SIDM, where collisional interactions dominate. Previous simulations at Milky Way scales Yang et al. (2025a) demonstrated that this mechanism reshapes inner halo profiles under strong scattering. Here, we extend this framework to simulations across dwarf to cluster scales, developing a conditioned parametric model to accurately capture gravothermal evolution near the resolution limit. We show that even modest inter-species interactions can induce mass segregation, increasing central densities while producing cores in dwarf halos. Crucially, unlike in one-component SIDM, the mass enclosed within the Einstein radius can also grow, amplifying strong lensing signals. Our results point to a unified scenario in which both cored dwarf galaxies and cuspy cluster substructures emerge naturally from SIDM with mass segregation.
II A two-component SIDM model with mass segregation
@@
Figure 1: Effective cross section per mass $\sigma_{\rm eff}/m$ as a function of $V_{\rm max}$ for various SIDM models. The self-interaction among the heavier component ($\chi_{H}-\chi_{H}$) and the inter-species interaction ($\chi_{H}-\chi_{L}$) in the SIDM2v model are represented by red and magenta solid curves. This model is featured by its capability of reproducing SIDM effects analogous to one dark matter component with $\sigma/m=0.3~{\rm cm^{2}/g}$ (SIDM03) in massive dwarf halos of mass around $M_{h}=10^{11}~\rm M_{\odot}$, as favored by the findings of Ref. Zhang et al. (2025b). The two orange stars feature the two kinds of strengths in the two-component SIDM2c model, which takes constant effective cross sections as in SIDM2v in a $M_{h}=10^{11}~\rm M_{\odot}$ halo with median concentration. For $V_{\rm max}\gtrsim 200~\mathrm{km/s}$, the inter-species interaction dominates over the intra-species interactions and remains significant up to the cluster scale at $V_{\rm max}\approx 1600~\mathrm{km/s}$. In clusters, the coexistence of both types of scattering yields density profiles consistent with observations. Particularly, the intra-species interaction remains well below $0.1~\rm cm^{2}/g$, as indicated by the light-red arrow.
@@
Two-component dark matter models naturally emerge in many particle models, such as those that mirror the asymmetries of the baryonic sector, where protons and electrons are stable and interact with each other Kaplan et al. (2009); Bansal et al. (2023); Petraki et al. (2012); Yang et al. (2025a). To incorporate mass segregation, we consider a two-component SIDM model in which the heavier species, $\chi_{H}$, has mass $m_{H}=3m_{L}$, with $\chi_{L}$ denoting the lighter component. We assume equal number densities and identical initial spatial distributions for both species. We introduce a model, SIDM2v, featuring velocity-dependent interactions for both identical (intra-species) and distinguishable (inter-species) scatterings. While both cross sections decrease with velocity, the inter-species interaction dominates at high velocities ($V_{\rm max}\gtrsim 200~\mathrm{km/s}$). Specifically, we adopt the Møller and Rutherford parametrizations for identical and distinguishable particle scatterings, following equations in Refs. Yang et al. (2025a); Yang and Yu (2022). The intra-species interaction is characterized by $\sigma_{0}/m_{H}=6.89~\rm cm^{2}/g$ and a velocity scale $w=275~\rm km/s$ (denoted $w_{H}$), while the inter-species interaction uses $\sigma_{0}/m_{H}=1.125~\rm cm^{2}/g$ and $w=2200~\rm km/s$. We assume that the same particle mediates all intra-species interactions, so that the $\chi_{L}-\chi_{L}$ interaction shares the same $\sigma_{0}$, with a velocity scale $w=(m_{H}/m_{L})w_{H}$.
@@
Figure 1 illustrates our model in terms of effective cross sections, defined as velocity- and angle-averaged differential cross sections weighted by a $v^{5}\sin^{2}\theta$ kernel and convolved with the dark matter velocity distribution. The $x$-axis corresponds to halo velocity scales via their maximum circular velocity, $V_{\rm max}$. The two orange stars indicate the intra- and inter-species interactions in a SIDM2c model without velocity dependence. Their strengths are set by matching the effective cross sections of SIDM2v in a benchmark halo of mass $10^{11}~\rm M_{\odot}$, yielding $\sigma_{0}/m=2.25~\rm cm^{2}/g$ for intra-species and $\sigma_{0}/m=1.125~\rm cm^{2}/g$ for inter-species interactions. This setup leads to time-dependent core formation in most dwarf halos of mass $\sim 10^{11}~\rm M_{\odot}$, comparable to one-component $\sigma/m=0.3~\rm cm^{2}/g$, as motivated by Ref. Zhang et al. (2025b).
@@
SIDM2c $\chi_{H}{\texttt{-}}\chi_{H}$ $\frac{\sigma_{H}}{m_{H}}=4.5\rm\frac{cm^{2}}{g}$
$\chi_{L}{\texttt{-}}\chi_{L}$ $\frac{\sigma_{L}}{m_{L}}=\frac{1}{3}\frac{\sigma_{H}}{m_{H}}$
$\chi_{H}{\texttt{-}}\chi_{L}$ $\frac{\sigma_{x}}{m_{H}}=\frac{1}{4}\frac{\sigma_{H}}{m_{H}}$
SIDM2v $\chi_{H}{\texttt{-}}\chi_{H}$ $\frac{\sigma_{H}}{m_{H}}=6.89\rm\frac{cm^{2}}{g}$ $w_{1}=275\ \rm\frac{km}{s}$
$\chi_{L}{\texttt{-}}\chi_{L}$ $\frac{\sigma_{L}}{m_{L}}=\frac{1}{3}\frac{\sigma_{H}}{m_{H}}$ $w_{2}=3w_{1}$
$\chi_{H}{\texttt{-}}\chi_{L}$ $\frac{\sigma_{x}}{m_{H}}=1.125\rm\frac{cm^{2}}{g}$ 2200 $\rm\frac{km}{s}$
SIDMx $\chi_{H}{\texttt{-}}\chi_{L}$ $\frac{\sigma_{x}}{m_{H}}=1.125\rm\frac{cm^{2}}{g}$
@@
Table 1: Benchmark models of two-component SIDM with mass segregation. The $\sigma_{0}/m$ and $w$ parameters are introduced to parametrize the scattering cross sections, following Ref. Yang et al. (2025a); Yang and Yu (2022) and considering Møller (Rutherford) scatterings for identical (distinguishable) particles. For identical particle scattering, a phase space suppression factor of $1/2$ is included to enable comparison with distinguishable inter-species scatterings, yielding the $\sigma_{0}/m=4.5/2{\rm cm^{2}/g}=2.25~\rm cm^{2}/g$ in the SIDM2c case.
III Numerical methods and modeling
@@
To accurately model signatures including baryonic and environmental effects, we employ controlled and cosmological simulations, combined with a parametric model for SIDM2c. The latter is constructed to obtain accurate halo profiles beyond the simulation resolution limit.
@@
Our simulations are based on the SIDM module developed on top of the Gadget-2 program Springel (2005); Springel et al. (2001) in Refs. Yang and Yu (2022); Yang et al. (2023), extended to handle two dark matter species with unequal masses. This module has been validated in both isolated and cosmological contexts. To establish two-component initial conditions without altering the system’s dynamical properties, we randomly select half of the simulation particles to be 50% more massive, and the other half to be 50% less massive, preserving the total system mass. For isolated halos, we generate Navarro-Frenk-White (NFW) profiles Navarro et al. (1997) in hydrostatic equilibrium using SpherIC Garrison-Kimmel et al. (2013), simulating with $N=10^{6}$ particles per run. The softening length is chosen to be a few $R_{\rm vir}/\sqrt{N}$ Power et al. (2003), where $R_{\rm vir}$ is the virial radius. A stellar component, modeled by the Hernquist profile Hernquist (1990), is included in the study with baryons. They are simulated as collisionless particles with the same mass as the dark matter particles. In the Supplemental Material, we provide a summary of the controlled simulations in a table. For cosmological simulations, we initialize a periodic box of size $400~{\rm Mpc}/h$, and perform zoom-in simulations of a $20~{\rm Mpc}/h$ region centered on a host halo of mass $10^{15}~{\rm M_{\odot}}/h$, with a softening length of one comoving $kpc/h$, where $h$ is the dimensionless Hubble parameter. The highest-resolution particles have average mass $m_{0}\equiv(m_{1}+m_{2})/2\approx 7\times 10^{7}~{\rm M_{\odot}}/h$. As shown in Ref. Yang et al. (2025a), mass segregation is negligible in CDM when particle masses are realistic, so we simulate CDM using a single-component model. We also apply a modified version of the Rockstar halo finder Behroozi et al. (2013) to accommodate two-component dark matter, a method validated and applied in Ref. Yang et al. (2025a). Using the resulting halo catalogs, we then build halo merger trees with the consistent-trees program Behroozi et al. (2013). For one-component SIDM in isolated halos, we employ the parametric framework of Ref. Yang et al. (2024), which has been extensively tested across controlled and cosmological settings Yang et al. (2025b).
@@
$\rho_{H}\nu_{H}^{2}\left(\frac{D}{Dt}\right)\ln\frac{\nu_{H}^{3}}{\rho_{H}}$ $=$ $-\frac{1}{4\pi r^{2}}\frac{\partial L_{H}}{\partial r}-R,$ (1)
$\rho_{H}\nu_{H}^{2}\left(\frac{D}{Dt}\right)\ln\frac{\nu_{L}^{3}}{\rho_{L}}$ $=$ $-\frac{1}{4\pi r^{2}}\frac{\partial L_{L}}{\partial r}+R,$
@@
Figure 2: Rescaled halo density profiles at 60% of the core collapse time $\tau\equiv t/t_{c}=0.6$, compared to a single-component (SIDM1c) model at the same evolution stage. The profiles of three halos of mass $2\times 10^{10}~M_{\odot}$ (solid) and three of mass $10^{11}~M_{\odot}$ (dashed) with median, $1\sigma$, and $1\sigma$ concentrations have almost identical shapes, illustrating the high quality of universality supporting the construction of the parametric model.
Figure 3: Dwarf halo evolution and core formation in one- and two-component SIDM. Left: Density profiles for the median halo with a collisionless stellar component at $t=0$, 33, 66, and 99~Gyr, with deeper colors corresponding to later times. Baryons (dotted tomato lines) induce higher DM densities (solid) compared to DM-only cases (dashed). Unlike DM-only evolution, central density increases even as the cores grow in size. The parametric model predictions are shown in dash-dotted lines. Right: Evolution of core size ($r_{c}$) for the median halo with (solid and dashed) and without (dotted) baryons. The core growth in SIDM2c remains comparable to that in SIDM03 in the presence of baryons.
In the long-mean-free-path regime, the SIDM effect scales with the scattering rate, hence $L_{H},L_{R}$ and $R$ are all proportional to the scattering cross section. In two-component models, the normalization of these quantities, parameterized by $\sigma_{s}/m$, can be absorbed into the evolution time, defining a rescaled evolution variable $t\sigma_{s}/m$. It follows that SIDM halos exhibit a universal evolutionary track when the mass ratio and the relative inter- to intra-species cross sections are fixed.
Figure 2 illustrates the accuracy of this conditioned universality by showing that the rescaled halo density profiles have almost the same shapes if evolved under SIDM2c for 60% of their core collapse times. We consider two halos of masses $2\times 10^{10}~\rm M_{\odot}$ and $10^{11}~\rm M_{\odot}$, with varied concentrations computed following the $c_{200}-M$ relation from Ref. Dutton and Macciò (2014) and assuming a logarithmic scatter of $0.11$ dex. Compared with the density profile of a halo in one-component SIDM (SIDM 1c) at the same gravothermal phase ($\tau\equiv t/t_{c}=0.6$, shown in dotted, tomato-colored curve), the profiles in SIDM2c have cored profiles and significantly higher inner densities.
@@
Figure 4: Distribution of the gravothermal phase $\tau$ from parametric model predictions for the SIDM2c model. The sample includes halos of masses higher than $10^{10}~\rm M_{\odot}$ and within $5~\mathrm{Mpc}/h$ of the host cluster, separated into cluster subhalos (Group A, light red) within the cluster’s virial radius and isolated halos (Group B, light blue). Means and standard deviations for both groups are shown in the inset. Most halos in both groups have cored inner profiles ($\tau<0.6$). Cluster subhalos, however, show a greater frequency of core collapsed systems ($\tau>1$), as a consequence of tidal stripping.
IV Dwarf scale cores
@@
We simulate the median-concentration halo with and without a collisionless stellar component, which follows median stellar-to-halo mass and size-mass relations. Figure 3 (left) presents results in the same style as Fig. 2 (right). While the halo densities in the baryonic cases (solid) are systematically higher than the DM-only cases (dashed), we found two crucial differences. One is that the stellar component develop cores in time, a feature directly addresses the observations. The other is that halos also have growing cores alongside with the increasing inner densities. This feature is precisely the opposite of the dark matter only case, where the core formation is accompanied with a decreasing inner halo density.
To track core growth clearly, we fit the halo density profiles with a $\beta 4$ model, defined as $\rho_{\beta 4}=\rho_{s}/((r/r_{s})^{4}+(r/r_{c})^{4})^{1/4}/(1+r/r_{s})^{2}$ Gilman et al. (2021); Yang et al. (2024), and plot the resulting core size $r_{c}$ evolution in Fig. 3 (right). After including baryons, the core size is significantly reduced due to gravitational contraction. Crucially, the core growth in SIDM2c appears close to that of the one-component SIDM03 after incorporating baryons.
@@
Figure 6: Comparison of simulated gravitational lenses in CDM (left), Model-2v (middle), and Model-x (right). Top panels show tangential (opaque) and radial (40% transparent) critical curves overlaid on projected density maps; bottom panels show corresponding caustics. The two-component models produce significantly more secondary caustics due to dense substructures. The GGSL cross section ($\sigma_{\rm GGSL}$) is computed summing over the areas enclosed by the secondary caustics whose effective Einstein radii are larger than $0.5^{\prime\prime}$ and residing within the $(-100^{\prime\prime},100^{\prime\prime})$ square region centered on the host halo.
V Strong lensing perturbers
We have seen that mass segregation can produce higher inner halo densities than one-component SIDM, even during core formation. A denser inner halo also accelerates the halo’s gravothermal evolution. This is intuitive from the dependence of the core-collapse time $t_{c}$ on halo concentration $c$. A dimensional analysis of $t_{c}$, translating the NFW parameters into mass and concentration, yields $t_{c}\propto c^{(a-7)/2}$ for a generic $v^{-a}$ cross section in one-component SIDM. For both a velocity-independent case ($a=0$) and the commonly used Coulomb-like scattering ($a=4$), $t_{c}$ decreases sharply as $c$ grows. Hence, a more compact halo is expected to have a shorter $t_{c}$.
Quantitatively, we calibrate the core collapse time of halos simulated in the SIDM2c model and find that it is close to that of a one-component SIDM model with $\sigma/m\approx 14~\text{cm}^{2}/\text{g}$, rather than the $0.3~\text{cm}^{2}/\text{g}$ value discussed earlier. As shown in Fig. 4, a small population of high concentration dwarf halos can evolve into dense, collapsing configurations under SIDM2c, explaining the observed dense dark substructures in strong lensing systems, commonly referred to as strong lensing perturbers Vegetti and Koopmans (2009); Vegetti et al. (2010, 2012); Hezaveh et al. (2016); Minor et al. (2021); Tajalli et al. (2025); Kong et al. (2025); Powell et al. (2025).

### v3

@@ Abstract
In two-component self-interacting dark matter (SIDM) models with inter-species interactions, mass segregation arises naturally from collisional relaxation, enhancing central densities and gravothermal evolution. We demonstrate that models with velocity-dependent interactions, both within and between species, can connect several small-scale observations while remaining consistent with cluster-scale constraints. This combination enables core formation in dwarf halos, where the presence of baryons increases the inner densities and enhances the predicted strong lensing signatures. Using cosmological and controlled simulations alongside an accurate parametric model, we present proof-of-principle examples showing that this framework can explain the structure of dark perturbers observed in strong lensing systems, and can enhance the efficiency of small-scale lenses by a factor of a few, in line with the excess reported in galaxy–galaxy strong-lensing observations. Importantly, mass segregation can enhance the Einstein radii of SIDM halos relative to their cold dark matter (CDM) counterparts, overcoming a key challenge in one-component SIDM scenarios. Our results present mass segregation in two-component SIDM as a self-consistent, testable framework with the potential to address multiple small-scale challenges in structure formation.
Self-interacting dark matter, Mass segregation, Strong lensing, N-body simulation, Parametric model
I 1. Introduction
@@
In this work, we demonstrate that SIDM with mass segregation provides a natural unified framework for resolving these tensions. Mass segregation in multi-component SIDM models arises as a consequence of collisions among particles of different masses. In the two-component case, inter-species scatterings lead to energy transfer from the heavier to the lighter species, resulting in denser inner halos Yang et al. (2025a); Patil and Fischer (2025). This process is negligible in CDM due to inefficient gravitational relaxation among particles (unlike stars Zhong and Shapiro (2025)), but becomes significant in SIDM, where collisional interactions dominate. Previous simulations at Milky Way scales Yang et al. (2025a) demonstrated that this mechanism reshapes inner halo profiles under strong scattering. Here, we extend this framework to simulations across dwarf to cluster scales, developing a conditioned parametric model to accurately capture gravothermal evolution near the resolution limit. We show that even modest inter-species interactions can induce mass segregation, increasing central densities while producing cores in dwarf halos. Crucially, unlike in one-component SIDM, the mass enclosed within the Einstein radius can also grow, amplifying strong lensing signals. Our results point to a unified scenario in which both cored dwarf galaxies and cuspy cluster substructures emerge naturally from SIDM with mass segregation.
II 2. A two-component SIDM model with mass segregation
@@
Fig. 1: Effective cross section per mass $\sigma_{\rm eff}/m$ as a function of $V_{\rm max}$ for various SIDM models. The self-interaction among the heavier component ($\chi_{\rm H}-\chi_{\rm H}$) and the inter-species interaction ($\chi_{\rm H}-\chi_{\rm L}$) in the SIDM2v model are represented by red and magenta solid curves. This model is featured by its capability of reproducing SIDM effects analogous to one dark matter component with $\sigma/m=0.3~{\rm cm^{2}/g}$ (SIDM03) in massive dwarf halos of mass around $M_{\rm h}=10^{11}~\rm M_{\odot}$, as favored by the findings of Ref. Zhang et al. (2025b). The two orange stars feature the two kinds of strengths in the two-component SIDM2c model, which takes constant effective cross sections as in SIDM2v in a $M_{\rm h}=10^{11}~\rm M_{\odot}$ halo with median concentration. For $V_{\rm max}\gtrsim 200~\mathrm{km/s}$, the inter-species interaction dominates over the intra-species interactions and remains significant up to the cluster scale at $V_{\rm max}\approx 1600~\mathrm{km/s}$. In clusters, the coexistence of both types of scattering yields density profiles consistent with observations. Particularly, the intra-species interaction remains well below $0.1~\rm cm^{2}/g$, as indicated by the light-red arrow.
@@
Two-component dark matter models naturally emerge in many particle models, such as those that mirror the asymmetries of the baryonic sector, where protons and electrons are stable and interact with each other Kaplan et al. (2009); Bansal et al. (2023); Petraki et al. (2012); Yang et al. (2025a). To incorporate mass segregation, we consider a two-component SIDM model in which the heavier species, $\chi_{\rm H}$, has mass $m_{\rm H}=3m_{\rm L}$, with $\chi_{\rm L}$ denoting the lighter component. We assume equal number densities and identical initial spatial distributions for both species. We introduce a model, SIDM2v, featuring velocity-dependent interactions for both identical (intra-species) and distinguishable (inter-species) scatterings. While both cross sections decrease with velocity, the inter-species interaction dominates at high velocities ($V_{\rm max}\gtrsim 200~\mathrm{km/s}$). Specifically, we adopt the Møller and Rutherford parametrizations for identical and distinguishable particle scatterings, following equations in Refs. Yang et al. (2025a); Yang and Yu (2022). The intra-species interaction is characterized by $\sigma_{0}/m_{\rm H}=6.89~\rm cm^{2}/g$ and a velocity scale $w=275~\rm km/s$ (denoted $w_{\rm H}$), while the inter-species interaction uses $\sigma_{0}/m_{\rm H}=1.125~\rm cm^{2}/g$ and $w=2200~\rm km/s$. We assume that the same particle mediates all intra-species interactions, so that the $\chi_{\rm L}-\chi_{\rm L}$ interaction shares the same $\sigma_{0}$, with a velocity scale $w=(m_{\rm H}/m_{\rm L})w_{\rm H}$.
@@
Fig. 1 illustrates our model in terms of effective cross sections, defined as velocity- and angle-averaged differential cross sections weighted by a $v^{5}\sin^{2}\theta$ kernel and convolved with the dark matter velocity distribution. The $x$-axis corresponds to halo velocity scales via their maximum circular velocity, $V_{\rm max}$. The two orange stars indicate the intra- and inter-species interactions in a SIDM2c model without velocity dependence. Their strengths are set by matching the effective cross sections of SIDM2v in a benchmark halo of mass $10^{11}~\rm M_{\odot}$, yielding $\sigma_{0}/m=2.25~\rm cm^{2}/g$ for intra-species and $\sigma_{0}/m=1.125~\rm cm^{2}/g$ for inter-species interactions. This setup leads to time-dependent core formation in most dwarf halos of mass $\sim 10^{11}~\rm M_{\odot}$, comparable to one-component $\sigma/m=0.3~\rm cm^{2}/g$, as motivated by Ref. Zhang et al. (2025b).
@@
Table 1: Benchmark models of two-component SIDM with mass segregation.
@@
SIDM2c $\chi_{\rm H}{\texttt{-}}\chi_{\rm H}$ $\frac{\sigma_{\rm H}}{m_{\rm H}}=4.5\rm\frac{cm^{2}}{g}$
$\chi_{\rm L}{\texttt{-}}\chi_{\rm L}$ $\frac{\sigma_{\rm L}}{m_{\rm L}}=\frac{1}{3}\frac{\sigma_{\rm H}}{m_{\rm H}}$
$\chi_{\rm H}{\texttt{-}}\chi_{\rm L}$ $\frac{\sigma_{\rm x}}{m_{\rm H}}=\frac{1}{4}\frac{\sigma_{\rm H}}{m_{\rm H}}$
SIDM2v $\chi_{\rm H}{\texttt{-}}\chi_{\rm H}$ $\frac{\sigma_{\rm H}}{m_{\rm H}}=6.89\rm\frac{cm^{2}}{g}$ $w_{1}=275\ \rm\frac{km}{s}$
$\chi_{\rm L}{\texttt{-}}\chi_{\rm L}$ $\frac{\sigma_{\rm L}}{m_{\rm L}}=\frac{1}{3}\frac{\sigma_{\rm H}}{m_{\rm H}}$ $w_{2}=3w_{1}$
$\chi_{\rm H}{\texttt{-}}\chi_{\rm L}$ $\frac{\sigma_{\rm x}}{m_{\rm H}}=1.125\rm\frac{cm^{2}}{g}$ 2200 $\rm\frac{km}{s}$
SIDMx $\chi_{\rm H}{\texttt{-}}\chi_{\rm L}$ $\frac{\sigma_{\rm x}}{m_{\rm H}}=1.125\rm\frac{cm^{2}}{g}$
@@
The $\sigma_{0}/m$ and $w$ parameters are introduced to parametrize the scattering cross sections, following Ref. Yang et al. (2025a); Yang and Yu (2022) and considering Møller (Rutherford) scatterings for identical (distinguishable) particles. For identical particle scattering, a phase space suppression factor of $1/2$ is included to enable comparison with distinguishable inter-species scatterings, yielding the $\sigma_{0}/m=4.5/2{\rm cm^{2}/g}=2.25~\rm cm^{2}/g$ in the SIDM2c case.
III 3. Numerical methods and modeling
@@
To accurately model signatures including baryonic and environmental effects, we employ controlled and cosmological simulations, combined with a parametric model for SIDM2c. The latter is constructed to obtain accurate halo profiles beyond the simulation resolution limit. We use the parametric model to mitigate resolution limitations in mapping CDM merger histories to SIDM2c profiles, and we adopt a conservative approach to the resulting lensing inferences.
@@
Our simulations are based on the SIDM module developed on top of the Gadget-2 program Springel (2005); Springel et al. (2001) in Refs. Yang and Yu (2022); Yang et al. (2023), extended to handle two dark matter species with unequal masses. This module has been validated in both isolated and cosmological contexts. To establish two-component initial conditions without altering the system’s dynamical properties, we randomly select half of the simulation particles to be 50% more massive, and the other half to be 50% less massive, preserving the total system mass. For isolated halos, we generate Navarro-Frenk-White (NFW) profiles Navarro et al. (1997) in hydrostatic equilibrium using SpherIC Garrison-Kimmel et al. (2013), simulating with $N=10^{6}$ particles per run. The softening length is chosen to be a few $R_{\rm vir}/\sqrt{N}$ Power et al. (2003), where $R_{\rm vir}$ is the virial radius. A stellar component, modeled by the Hernquist profile Hernquist (1990), is included in the study with baryons. They are simulated as collisionless particles with the same mass as the dark matter particles. In the Supplemental Material, we provide a summary of the controlled simulations in a table. For cosmological simulations, we initialize a periodic box of size $400~{\rm Mpc}/h$, and perform zoom-in simulations of a $20~{\rm Mpc}/h$ region centered on a host halo of mass $10^{15}~{\rm M_{\odot}}/h$, with a softening length of one comoving ${\rm kpc}/h$, where $h$ is the dimensionless Hubble parameter. The highest-resolution particles have average mass $m_{0}\equiv(m_{1}+m_{2})/2\approx 7\times 10^{7}~{\rm M_{\odot}}/h$. As shown in Ref. Yang et al. (2025a), mass segregation is negligible in CDM when particle masses are realistic, so we simulate CDM using a single-component model. We also apply a modified version of the Rockstar halo finder Behroozi et al. (2013) to accommodate two-component dark matter, a method validated and applied in Ref. Yang et al. (2025a). Using the resulting halo catalogs, we then build halo merger trees with the consistent-trees program Behroozi et al. (2013). For one-component SIDM in isolated halos, we employ the parametric framework of Ref. Yang et al. (2024), which has been extensively tested across controlled and cosmological settings Yang et al. (2025b).
@@
$\rho_{\rm H}\nu_{\rm H}^{2}\left(\frac{D}{Dt}\right)\ln\frac{\nu_{\rm H}^{3}}{\rho_{\rm H}}$ $=$ $-\frac{1}{4\pi r^{2}}\frac{\partial L_{\rm H}}{\partial r}-R,$ (1)
$\rho_{\rm L}\nu_{\rm L}^{2}\left(\frac{D}{Dt}\right)\ln\frac{\nu_{\rm L}^{3}}{\rho_{\rm L}}$ $=$ $-\frac{1}{4\pi r^{2}}\frac{\partial L_{\rm L}}{\partial r}+R,$
@@
Fig. 2: Rescaled halo density profiles at 60% of the core collapse time $\tau\equiv t/t_{\rm c}=0.6$, compared to a single-component (SIDM1c) model at the same evolution stage. The profiles of three halos of mass $2\times 10^{10}~\rm M_{\odot}$ (solid) and three of mass $10^{11}~\rm M_{\odot}$ (dashed) with median, $1\sigma$, and $1\sigma$ concentrations have almost identical shapes, illustrating the high quality of universality supporting the construction of the parametric model.
[…]
Fig. 3: Dwarf halo evolution and core formation in one- and two-component SIDM. (a) Density profiles for the median halo with a collisionless stellar component at $t=0$, 33, 66, and 99~Gyr, with deeper colors corresponding to later times. Baryons (dotted tomato lines) induce higher DM densities (solid) compared to DM-only cases (dashed). Unlike DM-only evolution, central density increases even as the cores grow in size. The parametric model predictions are shown in dash-dotted lines. (b) Evolution of core size ($r_{c}$) for the median halo with (solid and dashed) and without (dotted) baryons. The core growth in SIDM2c remains comparable to that in SIDM03 in the presence of baryons.
In the long-mean-free-path regime, the SIDM effect scales with the scattering rate, hence $L_{\rm H},L_{\rm L}$ and $R$ are all proportional to the scattering cross section. In two-component models, the normalization of these quantities, parameterized by $\sigma_{s}/m$, can be absorbed into the evolution time, defining a rescaled evolution variable $t\sigma_{s}/m$. It follows that SIDM halos exhibit a universal evolutionary track when the mass ratio and the relative inter- to intra-species cross sections are fixed.
Fig. 2 illustrates the accuracy of this conditioned universality by showing that the rescaled halo density profiles have almost the same shapes if evolved under SIDM2c for 60% of their core collapse times. We consider two halos of masses $2\times 10^{10}~\rm M_{\odot}$ and $10^{11}~\rm M_{\odot}$, with varied concentrations computed following the $c_{200}-M$ relation from Ref. Dutton and Macciò (2014) and assuming a logarithmic scatter of $0.11$ dex. Compared with the density profile of a halo in one-component SIDM (SIDM 1c) at the same gravothermal phase ($\tau\equiv t/t_{\rm c}=0.6$, shown in dotted, tomato-colored curve), the profiles in SIDM2c have cored profiles and significantly higher inner densities.
@@
Fig. 4: Distribution of the gravothermal phase $\tau$ from parametric model predictions for the SIDM2c model. The sample includes halos of masses higher than $10^{10}~\rm M_{\odot}$ and within $5~\mathrm{Mpc}/h$ of the host cluster, separated into cluster subhalos (group A, light red) within the cluster’s virial radius and isolated halos (group B, light blue). Means and standard deviations for both groups are shown in the inset. Most halos in both groups have cored inner profiles ($\tau<0.6$). Cluster subhalos, however, show a greater frequency of core collapsed systems ($\tau>1$), as a consequence of tidal stripping.
IV 4. Dwarf scale cores
@@
We simulate the median-concentration halo with and without a collisionless stellar component, which follows median stellar-to-halo mass and size-mass relations. Fig. 3a presents density profiles for the median halo with and without this stellar component. While the halo densities in the baryonic cases (solid) are systematically higher than the DM-only cases (dashed), we found two crucial differences. One is that the stellar component develops cores in time, a feature directly addresses the observations. The other is that halos also have growing cores along with the increasing inner densities. This feature is precisely the opposite of the dark matter only case, where the core formation is accompanied with a decreasing inner halo density.
To track core growth clearly, we fit the halo density profiles with a $\beta 4$ model, defined as $\rho_{\beta 4}=\rho_{\rm s}/((r/r_{\rm s})^{4}+(r/r_{c})^{4})^{1/4}/(1+r/r_{\rm s})^{2}$ Gilman et al. (2021); Yang et al. (2024), and plot the resulting core size $r_{c}$ evolution in Fig. 3b. After including baryons, the core size is significantly reduced due to gravitational contraction. Crucially, the core growth in SIDM2c appears close to that of the one-component SIDM03 after incorporating baryons.
@@
[…]
Fig. 6: Comparison of simulated gravitational lenses in CDM (a,d), Model-2v (b,e), and Model-x (c,f). Top panels show tangential (opaque) and radial (40% transparent) critical curves overlaid on projected density maps; bottom panels show corresponding caustics. The two-component models produce significantly more secondary caustics due to dense substructures. The GGSL cross section ($\sigma_{\rm GGSL}$) is computed summing over the areas enclosed by the secondary caustics whose effective Einstein radii are larger than $0.5^{\prime\prime}$ and reside within the $(-100^{\prime\prime},100^{\prime\prime})$ square region centered on the host halo.
V 5. Strong lensing perturbers
We have seen that mass segregation can produce higher inner halo densities than one-component SIDM, even during core formation. A denser inner halo also accelerates the halo’s gravothermal evolution. This is intuitive from the dependence of the core-collapse time $t_{\rm c}$ on halo concentration $c$. A dimensional analysis of $t_{\rm c}$, translating the NFW parameters into mass and concentration, yields $t_{\rm c}\propto c^{(a-7)/2}$ for a generic $v^{-a}$ cross section in one-component SIDM. For both a velocity-independent case ($a=0$) and the commonly used Coulomb-like scattering ($a=4$), $t_{\rm c}$ decreases sharply as $c$ grows. Hence, a more compact halo is expected to have a shorter $t_{\rm c}$.
Quantitatively, we calibrate the core collapse time of halos simulated in the SIDM2c model and find that it is close to that of a one-component SIDM model with $\sigma/m\approx 14~\text{cm}^{2}/\text{g}$, rather than the $0.3~\text{cm}^{2}/\text{g}$ value discussed earlier. As shown in Fig. 4, a small population of high concentration dwarf halos can evolve into dense, collapsing configurations under SIDM2c, explaining the observed dense dark substructures in strong lensing systems, commonly referred to as strong lensing perturbers Vegetti and Koopmans (2009); Vegetti et al. (2010, 2012); Hezaveh et al. (2016); Minor et al. (2021); Tajalli et al. (2025); Kong et al. (2025); He et al. (2025); Lei et al. (2025); Powell et al. (2025).